the mass of oxygen present in 0.5 mole of P4O10 is
1
Expert's answer
2019-11-21T06:26:30-0500

m(P4O10) = n(P4O10) • M(P4O10) = 0,5 mol • (31•4+16•10) g/mol = 0,5 • (136 + 160) = 0,5 • 296 = 148 g

Make a proportion:

Per 148 g of P4O10 - 296 g/mol

And per X g of Oxygen - 160 g/mol

So, X=148•160/296 = 80 g.


Answer: the mass of oxygen present in 0,5 mol of P4O10 is 80 grams.